OpenStudy (anonymous):

Simplify. square root 10 - 4 square root 10

OpenStudy (theviper):

\[\sqrt{10-4}\times \sqrt{10}\]\[\implies \sqrt{6}\times \sqrt{10}\]\[\implies \sqrt{60}\]

OpenStudy (anonymous):

the -4 isnt above the bracket and no multiply sign.. it's just \[\sqrt{10} -4 \sqrt{10}\]